# Model Editing at Scale leads to Gradual and Catastrophic Forgetting

### 15 Jan 2024 v1

@@ Abstract
Editing knowledge in large language models is an attractive capability to have which allows us to correct incorrectly learnt facts during pre-training, as well as update the model with an ever-growing list of new facts. While existing model editing techniques have shown promise, they are usually evaluated using metrics for reliability, specificity and generalization over one or few edits. We argue that for model editing to have practical utility, we must be able to make multiple edits to the same model. With this in mind, we evaluate the current model editing methods at scale, focusing on two state of the art methods - ROME and MEMIT. We find that as the model is edited sequentially with multiple facts, it continually forgets previously edited facts and the ability to perform downstream tasks. This forgetting happens in two phases - an initial gradual but progressive forgetting phase followed by abrupt or catastrophic forgetting phase. Both gradual and catastrophic forgetting limit the usefulness of model editing methods at scale - the former making model editing less effective as multiple edits are made to the model while the latter caps the scalability of such model editing methods. Our analysis also highlights other key limitations of ROME and MEMIT at scale. With our work, we push for the development and evaluation of model editing methods keeping scalability in mind.
@@ 1 Introduction
Editing knowledge in large language models (LLM) has recently emerged as a sought after capability for natural language processing (NLP) practitioners due to the huge retraining costs associated with LLMs. It is a known fact that LLMs memorize some part of the their training data (Radford et al., 2019; Carlini et al., 2022) in the form of stored knowledge. Model editing, sometimes also called knowledge editing [^1] In this paper, we will be using knowledge editing and model editing interchangeably, is the task of modifying existing knowledge or injecting new facts into the model Cohen et al. (2023), without updating the entire model. While knowledge in LLMs can be updated by continually pre-training models on new facts, this is not always viable due to the huge costs of training LLMs. Even then, LLMs are not perfect memorizers of training data Carlini et al. (2022) and will inevitably memorize facts incorrectly, thus requiring the capabilities of editing the incorrectly memorized facts. This is what makes model editing a very useful tool in the arsenal when working with LLMs.
@@
Various methods have been proposed over the years to do so. Some of these methods (De Cao et al., 2021; Mitchell et al., 2021) require training a hypernetwork Chauhan et al. (2023) that generates the new weights for the model being edited. Other methods (Meng et al., 2022a, b) directly update specific parts of the model after identifying location of stored facts inside it (Geva et al., 2020). The success of knowledge editing is usually evaluated along three dimensions (Yao et al., 2023; De Cao et al., 2021; Meng et al., 2022a, b) - (i) reliability, which measures if the post-edit model accurately recalls the newly edited fact, (ii) generalization, which measures if the post-edit model accurately recalls the newly edited facts for different phrasings of the edited fact, and (iii) locality (also known as specificity (Meng et al., 2022a, b)), which measures if the edited fact changes unrelated or neighboring facts.
While many of these methods have shown promise Yao et al. (2023), recent work analyzing the after-effects of these editing methods have highlighted the shortcomings of these methods. Specifically, while some of these editing methods rank high on reliability, generalization and locality metrics (Yao et al., 2023; Mitchell et al., 2021, 2022; Meng et al., 2022a, b), the edited knowledge is not used consistently by the model. Cohen et al. (2023) propose a new evaluation system where the "ripple effects" or implications of an edited fact are evaluated. An example of such ripple effects would be - if an edited fact updates the president of a country to the new president, then prompting for the birthplace of the president should output the birthplace of the new president. Li et al. (2023) extend this by introduce the concept of "knowledge conflict" and additional edit types like reverse-edits and round-edits, thus evaluating the logical consistency of model editing in more complex scenarios.
@@
Even though recent works are beginning to evaluate model editing more thoroughly, one very important aspect that is not given enough attention is making multiple edits on the same model. Most of prior works (De Cao et al., 2021; Mitchell et al., 2021; Meng et al., 2022a; Cohen et al., 2023; Li et al., 2023) focus on evaluating model editing performance when only a single edit is made to the model and find that making multiple edits, either in a sequential or batched manner, leads to significant degradation of performance (Mitchell et al., 2021, 2022; Meng et al., 2022b). In an ideal realization of editing knowledge in LLMs, these methods will be used to updates thousands if not hundreds of thousands of facts. When multiple edits were made to the same model sequentially, Meng et al. (2022b) see a significant decrease in performance for all methods, starting as low as 10 edits. This brings to attention a very important question - do these methods scale?
With the popularity of in-context learning Dong et al. (2022) and retrieval augmented generation (RAG) methods Lewis et al. (2020), knowledge editing only starts to become a viable alternative when the number of memories being edited or injected compete with the context length of the model. This means that knowledge editing has almost no practical use when used to edit just one fact and only starts to become useful at scale. Which is why in this paper, we keep scalability at the center when evaluating model editing in LLMs. We make multiple edits to the same model sequentially and analyse the effects of these edits on the model as a function of number of edits. Specifically, we use two best performing knowledge editing methods - ROME Meng et al. (2022a) and MEMIT Meng et al. (2022b) and make multiple edits on the same model. Along with the usual metrics for evaluating knowledge editing, we propose evaluating post-edit models on downstream tasks to make sure that model editing doesn’t hurt the regular functions of a model. We do that by evaluating the zero-shot performance of edited models on two downstream tasks from the GLUE benchmark Wang et al. (2018).
@@
We find that knowledge edits made by ROME and MEMIT are not as localized as they are believed to be. New edits consistently bleed into other facts stored in the model, including the previously edited facts. We also find the model editing methods are prone to gradual and catastrophic forgetting. To the best of our knowledge, our work is the first to associate model editing methods with catastrophic forgetting. We define gradual forgetting as a progressive loss of ability of the model to perform its regular functions as the model is continuously modified through knowledge edits. This includes forgetting the ability to recall previously edited facts as well doing downstream tasks. Gradual forgetting is followed by sudden or catastrophic forgetting (Goodfellow et al., 2013; Kirkpatrick et al., 2017), where the model gets crippled due to a single update made to the model. We call such edits made to the model disabling edits. A disabling edit decapitates the model - new knowledge edits are no longer successful on the model, the model forgets all previously edited facts and is unable to perform any downstream tasks. We also analyze different properties of and reasons behind such disabling edits. With this paper we highlight some serious limitations of current model editing techniques, especially their lack of robustness when scaled and call for further research in developing scalable model editing methods.
(a) Sample 1
(b) Sample 2
(c) Sample 3
@@
Figure 1: This figure shows the editing proficiency of ROME on GPT-2XL using three random samples of the CounterFact dataset. The dotted line represents the metric averaged over a past window size of 5, whereas the solid lines represent the metric averaged over a past window size of 50.
Specifically, we make the following contributions in this work:
  • We extensively evaluate ROME and MEMIT at scale and highlight some of their key features and limitations.
  • We propose evaluating model editing methods on downstream task performance in addition to knowledge editing metrics. This becomes even more important when editing facts at scale.
  • We find that models edited using ROME and MEMIT continuously undergo forgetting of previously edited facts as well as the ability to do downstream tasks. This forgetting takes place in two phases - a gradual but continuous phase followed by catastrophic forgetting. These behaviours limit the scalability of model editing. To the best of our knowledge, our work is the first to associate model editing methods with catastrophic forgetting.
  • We discover disabling edits - these are single knowledge edits that cause the model to lose all function. We find that disabling edits are fundamental to ROME, while MEMIT is more robust to such single disabling edits.
2 Dataset and Methods
In this paper, we evaluate two most popular and best performing knowledge editing methods in literature - Rank-One Memory Editing (ROME) Meng et al. (2022a) and Mass-Editing Memory in a Transformer (MEMIT) Meng et al. (2022b). We evaluate these models on the CounterFact dataset Meng et al. (2022a), which contains 21,919 counterfactual statements. Each datapoint in the dataset is a triplet of the type (subject, relation, object). The dataset is created such that the target object is less likely than the original object in the relation triplet. This makes each datapoint in the dataset much harder to edit. In this preprint, we perform all experiments on GPT-2XL Radford et al. (2019) due to computational constraints. These experiments will be expanded to other models in future iterations of the paper.
@@ 3 Scaling ROME
In this section, we evaluate the performance of Rank-One Memory Editing (ROME) method Meng et al. (2022a) when multiple edits are made to the same model. These edits are made using the CounterFact dataset. To perform these experiments, we create three random subsets of 1000 examples from the CounterFact dataset and sequentially edit the model on the selected datapoints. We do so to find patterns that are independent of the effect of the order in which facts are edited. In this paper, we present three representative subsets that represent three different ways in which model editing dynamics proceed for the ROME method. There is no knowledge conflict Li et al. (2023) in any of these samples as the same subject is not edited twice in a subset.
(a) Sample 1
(b) Sample 2
(c) Sample 3
Figure 2: This figure shows the percentage of facts forgotten as a function of number of edits made to GT-2XL using ROME on the same three samples as used in Figure 1. The percentage is calculated over the number of facts that were successfully edited up to given number of edits.
@@ 3.1 Editing Proficiency at Scale
Figure 1 shows the different metrics used for measuring editing proficiency of ROME on GPT-2XL as a function of the number of edits made to the model. The dotted lines represent average over a past window of size 5, which is the average of the given metric over 5 previous edits. The solid lines represents the average metric over a window size of 50. We find that the efficacy score, which measures the success of an edit, is almost 100% until a point where it begins to decline, sometimes very steeply. The three samples show that this point of steep decline can come as early as 100 edits, or as late as 800 edits made to the model  [^3] Experiments with other subsets have shown that this inflection point can occur at as early as 10 edits made to the model.. While ROME is successful at making edits to the model prior to this inflection point, its neighborhood accuracy still consistently declines as more edits are made to the model, indicating that the edits made start to bleed into other fact stored in the model, as more edits are made to the model. We will provide more evidence for this in later sections.
To summarize, we find that ROME is extremely proficient at successfully editing knowledge but only up to a certain point, after which, suddenly the method breaks down. Although even before that point is reached, the edited knowledge does begins to impact the unrelated knowledge stored in the model, as shown by the consistently declining neighborhood score.
@@ 3.2 Gradual and Catastrophic Forgetting
As new facts get added successfully to the model, is the model able to remember previously edited facts?
Figure 2 shows the number of previously correctly edited facts that get forgotten as a function of new edits made to the model. Firstly, we want to point the reader to clear spikes in all three graphs that occur at approximately the the same locations where the efficacy score in Figure 1 starts to decline. This indicates that the inflection point observed in Figure 1 has additional consequences other than just the loss of editability of the model..
Prior to this inflection point, Figure 2 shows an increasing relationship between the number of forgotten facts and the number of edits made to the model [^4] The graphs in figure 2 are evaluated after every 10 edits for computational reasons.. For example, for Sample 1, the model forgets approximately 30% of edited facts after 600 edits made to the model. Thus prior to the inflection point, we see a region where the model gradually forgets the previously edited facts. Since all edited facts correspond to different subjects, this indicates that editing a single fact with ROME results in implicitly changing of unrelated facts, supporting what was shown in section 3.1. After this region of gradual forgetting of facts, we reach an inflection point where we find that a catastrophically large number of facts are forgotten by the model. This is the same point where any further knowledge editing also suddenly starts to become unsuccessful (Figure 1).
(a) Sample 1
(b) Sample 2
(c) Sample 3
Figure 3: This figure shows the downstream performance of editing GPT2-XL using ROME on SST2 and MRPC tasks.
This phenomenon of sudden forgetting of a huge number of facts is a realization of catastrophic forgetting in machine learning literature (Goodfellow et al., 2013; Kirkpatrick et al., 2017). Catastrophic forgetting is defined as the sudden loss of ability of a model to perform a prior task when it is further trained to perform a new task. The phenomenon observed in the above example is a perfect realization of how "catastrophic" or abrupt catastrophic forgetting can be, where it literally "forgets" an exploding number of facts with one gradient update. To the best of our knowledge, our work is the first to show that knowledge editing methods are also prone to catastrophic forgetting. But is catastrophic forgetting just limited to abruptly forgetting previously edited facts? In the next section, we show that it goes beyond just forgetting of previously edited facts.

### 20 Feb 2024 v2

@@ Abstract
Editing knowledge in large language models is an attractive capability that allows us to correct incorrectly learned facts during pre-training, as well as update the model with an ever-growing list of new facts. While existing model editing techniques have shown promise, they are usually evaluated using metrics for reliability, specificity and generalization over one or few edits. We argue that for model editing to have practical utility, we must be able to make multiple edits to the same model. With this in mind, we evaluate current model editing methods at scale, focusing on two state of the art methods - ROME and MEMIT. With the lens of scalability, we evaluate model editing methods for three crucial properties - editing proficiency, fact forgetting and downstream performance. We find that as a model is edited sequentially with multiple facts, it continually becomes less editable, forgets previously edited facts and loses the ability to perform downstream tasks. For ROME and MEMIT, this "forgetting" happens in two phases - an initial gradual but progressive forgetting phase followed by an abrupt or catastrophic forgetting. Both gradual and catastrophic forgetting limit the usefulness of model editing methods at scale - the former makes model editing less effective as multiple edits are made to the model while the latter caps the scalability of such model editing methods. Our analysis also highlights other key limitations of ROME and MEMIT at scale. With our work, we push for better evaluation of model editing and development of model editing methods keeping scalability in mind.
@@ 1 Introduction
Editing knowledge in large language models (LLM) has recently emerged as a sought after capability for natural language processing (NLP) practitioners. It is a known fact that LLMs memorize some part of their training data (Radford et al., 2019; Carlini et al., 2022). Model editing, sometimes also called knowledge editing [^1] In this paper, we use knowledge editing and model editing interchangeably, is the task of modifying existing knowledge or injecting new facts into the model Cohen et al. (2023), without updating the entire model. While knowledge in LLMs can be updated by continually pre-training models on new facts, this is not always viable due to the huge costs of training LLMs and an ever-growing list of facts. Even then, LLMs do not perfectly memorize training data Carlini et al. (2022) and will inevitably memorize facts incorrectly, thus requiring the capabilities of model editing. This is what makes model editing a very useful tool in the arsenal when working with LLMs.
@@
Various methods have been proposed over the years to do so. Some of these methods (De Cao et al., 2021; Mitchell et al., 2021) require training a hypernetwork Chauhan et al. (2023) that generates new weights for the model being edited. Other methods (Meng et al., 2022a, b) directly update specific parts of the model after locating stored facts inside it (Geva et al., 2020). The success of knowledge editing is usually evaluated along three dimensions (Yao et al., 2023; Meng et al., 2022a, b) - (i) reliability, which measures if the post-edit model accurately recalls the newly edited fact, (ii) generalization, which measures if the post-edit model accurately recalls the newly edited facts for different phrasings of the edited fact, and (iii) locality (also known as specificity (Meng et al., 2022a, b)), which measures if the edited fact changes unrelated or neighboring facts.
@@
Most prior works (De Cao et al., 2021; Mitchell et al., 2021; Meng et al., 2022a; Cohen et al., 2023; Li et al., 2023b) focus on evaluating model editing performance by making one edit at a time. While some recent works (Mitchell et al., 2021, 2022; Meng et al., 2022b) have begun evaluating model editing with multiple edits made to the same model, we find the evaluation in this setting to not be exhaustive. In an ideal realization of model editing, these methods should be used to update thousands if not hundreds of thousands of facts. When multiple edits were made to the same model sequentially, Meng et al. (2022b) see a significant decrease in performance for all methods, starting as low as 10 edits. This brings to attention a very important question - do these methods scale? Which is why in this paper, we keep scalability at the center when evaluating model editing.
Dataset Name Query
zsRE
Who was the architect of Villa Kampen?
CounterFact
Porto is a twin city of
Table 1: Examples of input prompts from zsRE and CounterFact.
In this paper, we present a framework to study the dynamics of model editing at scale. We make multiple edits to the same model sequentially and analyze the effects of these edits on the model as a function of the number of edits. Specifically, we compare model editing at scale for three of the most popular model editing methods - MEND Mitchell et al. (2021), ROME Meng et al. (2022a) and MEMIT Meng et al. (2022b) against a fine-tuning baseline, and make multiple edits on the same model. Along with the usual metrics for evaluating model editing, we propose evaluating post-edit models for three important properties - editing proficiency, fact forgetting, and downstream task performance.
@@
We find that ROME and MEMIT outperform MEND and fine-tuning baselines at scale, but edits made by ROME and MEMIT are not as localized as previously believed to be. We show that new edits consistently bleed into other facts stored in the model. We also find the model editing methods are prone to gradual and catastrophic forgetting. To the best of our knowledge, our work is the first to associate model editing methods with catastrophic forgetting. We define gradual forgetting as a progressive loss of ability of the model to perform its regular functions as the model is continuously modified through knowledge edits. This includes forgetting the ability to recall previously edited facts as well doing downstream tasks. For ROME and MEMIT, gradual forgetting is followed by sudden or catastrophic forgetting (Goodfellow et al., 2013; Kirkpatrick et al., 2017), where the model gets crippled due to a single update made to the model. We name these edits - disabling edits. A disabling edit decapitates the model - new knowledge edits are no longer successful on the model, the model forgets all previously edited facts and is unable to perform any downstream tasks. We also study different properties of these disabling edits. With this paper we highlight some serious limitations of current model editing techniques, especially their lack of robustness when scaled and call for further research in developing scalable model editing methods. Our code can be found here.
2 Methods, Models and Datasets
In this paper, we focus on two prominent model editing methods in literature - Rank-One Memory Editing (ROME) Meng et al. (2022a) and Mass-Editing Memory in a Transformer (MEMIT) Meng et al. (2022b), whereas Model Editor Networks using Gradient Decomposition (MEND) Mitchell et al. (2021) and fine-tuning are used as baselines. MEND is a hypernetwork Chauhan et al. (2023) based model editing method that generates the weight updates for the model being edited. ROME and MEMIT first localize knowledge within a model using causal tracing Vig et al. (2020) and then update the weights of the selected layers to inject knowledge. The major difference between ROME and MEMIT is that while ROME works under the assumption that knowledge in LLMs can be updated by updating a single layer, MEMIT updates the weights of multiple layers. We evaluate these model editing methods over two models - GPT2-XL (1.5B) Radford et al. (2019) and GPT-J (6B) Wang and Komatsuzaki (2021). Note that all prior works (Mitchell et al., 2021; Meng et al., 2022a, b) edit base language models and not chat models. We refer the reader to appendix A.1 for a detailed survey of model editing techniques. This section will be added to the main paper upon acceptance.
Knowledge editing is usually evaluated on two datasets - the zsRE (zero-shot relation extraction) dataset Levy et al. (2017) and the CounterFact dataset Meng et al. (2022a). The zsRE dataset contains facts in the form of question-answer (QA) pairs created from Wikipedia. A key distinction between zsRE and CounterFact datasets is that zsRE contains true facts, which are easier for the model to learn, whereas CounterFact contains counterfactual examples where the new target has lower probability when compared to the original answer Meng et al. (2022a). Examples from these datasets can be seen in Table 1, which highlights another key difference between the two datasets. The queries in the zsRE dataset are present in the form of questions whereas for the CounterFact dataset, the queries are in the form of a prompt and is followed by an answer, which is a more natural formulation for base language models as they are trained to complete a sentence.
To check for the applicability of the zsRE dataset for editing base models, we first edit a fact in the QA format (as shown in Table 1), and then evaluate the success of the edit by prompting the question in a text completion format ("The architect of Villa Kampen is" ). We find that the model is unable to produce the correct answer in approximately 70% scenarios when prompted in a text completion format after successful edits in the QA format. This shows that facts edited using the zsRE dataset do not become a part of the text generation process (details in Appendix A.2). We thus choose the CounterFact dataset for our experiments.
The CounterFact dataset contains 21,919 counterfactual statements. Each datapoint in the dataset is a triplet of the type (subject, relation, object). The dataset is created such that the target object is less likely than the original object in the relation triplet. CounterFact is thus also a more difficult dataset Meng et al. (2022a) as we are going against the knowledge of the model, which is exactly what we do when we edit or inject facts in a model.
(a) FT-C
(b) MEND
(c) ROME
(d) % Facts Forgotten
@@ 2 Methods, Models and Datasets
Figure 1: This figure shows the editing proficiency of FT-C, MEND, and ROME on GPT-J (6B). The dotted line represents the metric averaged over a past window size of 5, whereas the solid lines represent the metric averaged over a past window size of 50. Figure 0(d) show the percentage of previously edited facts forgotten as a function of number of edits.
@@ 3 Scaling ROME
In this section, we evaluate the performance of Rank-One Memory Editing (ROME) method Meng et al. (2022a) when multiple sequential edits are made to the same model. We compare the performance of ROME with MEND and fine-tuning (FT-C) baseline. We use the standard implementation of ROME and MEND. For the fine-tuning baselines, we fine tune the same layer being edited by ROME and constrain the norm of change in weights. Our experiments show that if norm is not constrained during fine-tuning, its leads to immediate model degradation. We provide more elaborate implementation details in appendix A.3.
Edits to the models are made using the CounterFact dataset. To perform these experiments, we create four random subsets of 1000 examples from the CounterFact dataset and sequentially edit the model on the selected datapoints. We do so to find patterns that are independent of the effect of the order in which facts are edited. There is no knowledge conflict Li et al. (2023b) in any of these samples as the same subject is not edited twice.
@@ 3.1 Editing Proficiency at Scale
We first begin by evaluating editing proficiency of ROME as multiple edits are made to the model sequentially. This is done by analyzing the success of a new edit as a function of number edits made. To do so, we measure three metrics, following the convention of Meng et al. (2022a) - efficacy score, paraphrase score, and neighborhood score.
@@
Paraphrase score (PS) measures if the model is able to recall the edited fact with larger probability when prompted with a paraphrase of the sentence that was used to edit the fact. It is measured as true if $P(newfact)>P(oldfact)$ for a paraphrased prompt. Paraphrase score represents the generalization ability of the model editing method.
Neighborhood score (NS) measures the effect of editing the model on related facts with a different subject, and is measured true if $P(neighborhoodfact)>P(newfact)$. Neighborhood score represents specificity of the editing method, and is measured on a set of distinct but semantically related subjects. In this scenario, we want the neighborhood facts to not be affected by model editing.
Figure 1 shows the different metrics used for measuring editing proficiency of FT-C, MEND, and ROME on GPT-J (6B) as a function of the number of edits made to the model for one of the four samples. Additional experiments for other samples as well as GPT2-XL are shown in appendix A.5.1. The dotted lines represent average metric over a past window of size 5, which is the average of the given metric over 5 previous edits. The solid lines represents the average metric over a window size of 50.
We find that both FT-C and ROME are able to successfully edit facts sequentially, whereas MEND is unable to make multiple sequential edits to the same model, as shown by the efficacy score. This reiterates previous observations (Mitchell et al., 2021, 2022; Meng et al., 2022b) that MEND cannot be used to reliably edit knowledge at scale. For ROME, we find that the efficacy score, which measures the success of an edit, is almost 100% until a point where it begins to decline. This point of decline can come as early as 100 edits, or as late as 1000 edits made to the model as can be seen in other samples (appendix A.5.1). Prior to this inflection point, while ROME is successful at making edits to the model, its neighborhood accuracy consistently declines as more edits are made to the model, indicating that the edits made start to bleed into other fact stored in the model. We will provide more evidence for this in later sections. These trends are consistent across multiple samples and multiple models.
(a) FT-C
(b) MEND
(c) ROME
(d) Distance
Figure 2: This figure shows the downstream performance of editing GPT2-J on four GLUE tasks for different model editing methods. Figure 1(d) shows the the normalized distance between the edited layer and its original weights.
@@ 3.2 Gradual and Catastrophic Forgetting
As new facts get added successfully to the model, is the model able to remember previously edited facts? This is the question we try to answer in this section. Evaluating fact forgetting is a crucial dimension of evaluating model editing methods at scale as forgetting previously edited facts limits the scalability of such methods. Additionally, forgetting is a direct indication of locality. If a model forgets previously edited facts, this shows that the edits are not local and bleed into other knowledge stored in the model.
Figure 0(d) shows the number of previous correctly edited facts that get forgotten as a function of new edits made to GPT-J. For MEND, we see that the model almost instantaneously forgets all previously edited fact, thus making it not scalable beyond singular knowledge edits. For FT-C, we find that the model forgets previously edited facts rapidly as a function of newer edits made to the model, and at a time only retains a handful of prior edits. This also means that edits made using FT-C are highly non-local. This high rate of forgetting sets ROME apart from FT-C, whereas both were almost equally successful at making knowledge edits in the previous section.
For ROME, Figure 0(d) initially shows a slowly increasing relationship between the number of forgotten facts and the number of edits made to the model [^3] The graphs in figure 0(d) are evaluated after every 10 edits for computational reasons. at a rate which is much smaller than the forgetting rate of FT-C. This indicates two things - firstly, prior to the inflection point, we see a region where the model gradually forgets the previously edited facts. Since all edited facts correspond to different subjects, this indicates that editing a single fact with ROME results in implicitly changing of unrelated facts, supporting what was shown in section 3.1. Thus, edits made using ROME are not as local as previously believed to be. Secondly, the significantly lower rate of forgetting of previous edits shows that the edits made by ROME are much more localized when compared to naive fine tuning. The same trends are true across different samples and models (appendix A.5.2).
After this region of gradual forgetting of facts for ROME, we reach an inflection point where we find that a catastrophically large number of facts are forgotten by the model. This is the same point where any further knowledge editing also starts to slowly become unsuccessful using ROME (Figure 0(c)). This phenomenon of sudden forgetting of a huge number of facts is a realization of catastrophic forgetting in machine learning literature (Goodfellow et al., 2013; Kirkpatrick et al., 2017). Catastrophic forgetting is defined as the sudden loss of ability of a model to perform a prior task when it is further trained to perform a new task. The phenomenon observed in the above example is a perfect realization of how "catastrophic" or abrupt catastrophic forgetting can be, where it literally "forgets" an exploding number of facts with one gradient update. To the best of our knowledge, our work is the first to show that model editing methods are also prone to catastrophic forgetting.
But is catastrophic forgetting just limited to abruptly forgetting previously edited facts? In the next section, we show that it goes beyond that.